OpenStudy (anonymous):

Use Coulomb's law to calculate the ionization energy in kJ/mol of an atom composed of a proton and an electron separated by 256.00pm.

OpenStudy (anonymous):

For one particular atom, the potential energy is \[U = \frac{e^2}{4\pi\epsilon_0 R} \] So for a mole of atoms, \[U_{total} = \frac{N_A e^2}{4\pi\epsilon_0 R} \] where N_A is Avogadro's number. I calculate \[U_{total} \approx 2.39 \space kJ\]

OpenStudy (anonymous):

oops, make that 239